# Studying propagating turbulent structures in the near wake of a sphere using Hilbert proper orthogonal decomposition

### v1

@@ 4.4 Propagating Structures in the Wake of a Sphere
The contribution of the pairs of POD modes identified to the total TKE of $\boldsymbol{X}$ is given in table 6.
@@
In order to determine the nature and effects of the propagating structures identified from the POD modes, the paired POD modes were used to phase average the instantaneous velocity fluctuations
@@
where $i$ and $j$ are paired POD modes, and $N_{\phi}$ is the number of instantaneous velocity fields included for the phase angle $\phi$. The mean velocity was added to the phase-averaged velocity fluctuations to determine the phase-averaged velocity
@@
from which the phase-averaged vorticity is defined as
@@
The phase-averaged planar Reynolds stress is given by the phase average of the product of the streamwise and transverse components of the instantaneous velocity fluctuations. The TKE of the phase-averaged velocity is given by
@@
Figure 14: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the first and second POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
Figure 15: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the first HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
The motion represented by the first and second POD modes, shown in figure 14, and the first HPOD mode, shown in figure 15, is a flapping of the wake. The phase-averaged velocity from using the first and second POD modes and the first HPOD mode, shown in figure 14a and 15a, respectively, are qualitatively indistinguishable, as is the phase-averaged vorticity, shown in figure 14b and 15b, respectively. The structures in the planar Reynolds stresses of the phase-average of POD modes, shown in figure 14c, are longer and more intense for $x/D<3$ than those in the first HPOD mode, shown in figure 15c, and disappear after $x/D=3$. The structures in the planar Reynolds stress of the HPOD mode are shorter and more consistent in size across the measured domain. The intensity of the structures is lower than those in the POD modes, but decays more slowly, with structure visible past $x/D=3$. The phase-averaged TKE of the POD modes, shown in figure 14d, is concentrated in periodically shed packets which initially increase in intensity until $x/D=2$, and then decay by $x/D=3$. The TKE of the HPOD mode, shown in figure 15d, reaches exhibits the same periodic structures as the POD modes, however, the intensity is increased. This is consistent with the TKE of the analytic signal of the turbulent fluctuations being significantly higher than that of the actual turbulent fluctuations.
@@
Figure 16: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the fourth and fifth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
Figure 17: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the second HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
The structure represented by the fourth and fifth POD modes and the second HPOD mode is a pulsating in the streamwise fluctuations, as can be seen from the deformation of the low velocity region behind the sphere, shown in figure 16a and 17a, respectively. The streamwise pulsation can also be seen in the phase-averaged out-of-plane vorticity, with the high-vorticity regions either side of the sphere extend further toward the centreline close to the sphere in figures 16b(ii) and 17b(ii). This extension toward the centreline occurs around $x/D=2$ in figures 16b(v) and 17b(v). While the structures in the planar Reynolds stress are highly variable in shape and size in the phase-averaged POD modes, shown in figure 16c, they are consistent in size and shape in the HPOD mode, shown in figure 17c. The size and shape of the phase-averaged TKE of the POD modes, shown in figure 16d, is more consistent than that of the planar Reynolds stresses. These structures are similar to, but less planar symmetric, than the structures in the phase-averaged TKE of the HPOD mode, which also exhibit a higher intensity with a more consistent pattern. The intensity of the TKE increases from directly behind the sphere until $x/D=2$, as seen in figure 17d(ii), before decreasing further downstream, as seen in figure 17d(iv).
@@
Figure 18: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the sixth and eighth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
Figure 19: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the third HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
The sixth and eighth POD modes represent a flapping motion in the wake, similar to that of the first and second POD modes but with a shorter wavelengths. Thus, the phase-averaged velocity, shown in figure 18a, and phase-averaged vorticity, shown in figure 18b, are similar to those of the first and second POD modes. The smaller structures are evident in the phase-averaged planar Reynolds stress, shown in figure 18c, particularly near the sphere. Further downstream of the sphere, the structures increase in streamwise extent, most notable in figure 18c(iv). The smaller structures in the phase-averaged TKE, shown in figure 18d, are more definitively separated across the centreline of the sphere than in the case of the first and second POD modes, where the larger structures extended to the centreline. The structures in the phase-averaged planar Reynolds stresses of the third HPOD mode, shown in figure 19c, are more consistent in size, shape and arrangement than those of the POD modes. There is also a region with low planar Reynolds stress around $x/D=2$ for all phase angles shown in figure 19, which reflects the smaller structures in the streamwise turbulent fluctuations in this regions, shown in figure 7c. The structures in the phase-averaged TKE are also more clearly defined for the HPOD modes, as shown in figure 19d.
@@
Figure 20: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the ninth and tenth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
Figure 21: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the fourth HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
The structure represented by the ninth and tenth POD modes, and the fourth HPOD mode, is similar to that represented by the fourth and fifth POD modes, and the second HPOD mode, namely periodic streamwise fluctuations in the wake. The phase-averaged velocity, shown in figures 20a and 21a for the POD modes and HPOD mode, respectively, and the phase-averaged vorticity, shown in figures 20b and 21b for the POD modes and HPOD mode, respectively, are qualitatively similar to those of the lower order modes. The structures in the planar Reynolds stress phase-averaged using the POD modes, shown in figure 20c, are small and round in the near wake of the sphere and larger and more ovular further downstream, with the latter being angles away from the centreline in accordance with the V-shaped structures in the POD modes. While the arrangement of these structures is variable, particularly close to the sphere, in the phase-average using the POD modes, the planar Reynolds stress of the HPOD mode, shown in figure 21c, clearly shows the round structures for $x/D<2$ and more V-shaped structures for $x/D>2$. Both the POD modes and HPOD modes show a noticeable reduction in the planar Reynolds stress around $x/D=2$, which delineates the rounder structures from the V-shaped ones. However, the HPOD modes display this more prominently than the POD modes, which also exhibit additional structures in the very near wake which are filtered out in the HPOD modes. The change between round structures near the sphere to more V-shaped structures is reflected in the phase-averaged TKE, shown in figure 20d and 21d, for the POD modes and HPOD mode, respectively. This also shows that the structures close to the sphere are more intense and decay in intensity toward $x/D=2$, with the V-shaped structures increasing in intensity around $x/D=3$ before decaying in intensity downstream. As in the case of the planar Reynolds stress, the structures in the TKE are more clearly defined in the HPOD mode due to the filtering of the Hilbert transform.
@@
The structures elucidated by the POD mode pairs identified from the HPOD modes differ from those in the corresponding HPOD modes. The primary difference between these methods results from the filtering inherent in the Hilbert transform, as the non-propagating structures are given less importance in the HPOD, while propagating structures are amplified. The phase-averaging of the resulting modes thus results in more consistently shaped structures making it easier to identify and interpret the propagating structures. However, the filtering out of non-propagating structures may result in a poorer representation of the flow structure where the flow is three-dimensional. While applying the Hilbert transform directly to the POD modes is more computational efficient than performing the HPOD for pairing POD modes, it also forcibly propagates structures in the flow which may not propagate in reality. Thus, both methods of pairing POD modes introduce non-physical artefacts into the analysis of the flow and this should be considered when interpreting results.
@@ 5 Concluding Remarks
The velocity in the wake of a sphere at $Re_{D}=7780$ was measured using 2C-2D MCCDPIV, and POD and HPOD were applied to the turbulent fluctuations. The resulting POD and HPOD modes were compared to determine if the HPOD modes, which accentuate propagating structures, could be used to match POD modes which represented the same structure at different phase angles. The Hilbert transform of the POD modes was then used as a substitute for the HPOD modes in order to determine paired POD modes without performing the computationally-intensive HPOD. The POD mode pairs identified using these methods were then used to phase average the turbulent fluctuations to examine the propagating structures in the wake of the sphere.
The leading HPOD modes are shown to represent the same structures as the leading POD modes, with a greater emphasis on the propagating structures in the former. While each HPOD mode consists of a real and imaginary part representing a phase-shift of 90 in the mode, this comes at the cost of filtering from the Hilbert transform. As a result, it can be difficult to differentiate between what is part of the flow structures and what is an artefact of this filtering, particularly in the case of two-dimensional measurements of a three-dimensional flow. Performing the Hilbert transform on the POD modes to create an analytic signal of the POD modes is shown to identify the same pairs of POD modes identified using the HPOD modes. This method allows the same structures identified using the HPOD modes to be identified, without introducing the filtering resulting from the HPOD. The phase average of the turbulent fluctuations using the identified POD mode pairs represented flapping and pulsating motions in wake of the sphere.
As the velocity fields in the current work are not time-resolved, the Hilbert transform was performed in the streamwise direction, and thus the structures identified are those which propagate in that direction. As a result, the temporal information of the structures is not given by either decomposition. However, the paired POD modes gave insight into the dynamics of the structures in the turbulent fluctuations when the complexity of the flow made identifying related POD modes challenging.

### v2

@@ 4.4 Propagating Structures in the Wake of a Sphere
Table 6 compares the contributions of the identified POD mode pairs and their corresponding HPOD modes to the total turbulent kinetic energy (TKE). The first pair of POD modes (modes 1 and 2) represents the largest contribution to TKE, at 8.24%, which slightly exceeds the 7.70% contribution of the first HPOD mode. Subsequent pairs contribute progressively less energy, with the fourth pair (modes 9 and 10) accounting for 2.60% of the total TKE. The corresponding HPOD modes exhibit similar trends, although their contributions differ from those of the POD mode pairs. The first two HPOD modes contribute more proportionally to the analytic signal than the corresponding POD mode pairs contribute to the original data, whereas the third and fourth HPOD modes contribute less than their POD counterparts. This is consistent with the higher-order POD modes containing a greater proportion of non-propagating structures, which are excluded from the leading HPOD modes.
@@
In order to examine the nature and effects of the propagating structures identified from the HPOD and POD modes, the HPOD modes and corresponding POD mode pairs were employed to phase-average the instantaneous velocity fluctuations. The phase-averaged velocity fluctuations based on the phase angle of the $i^{th}$ and $j^{th}$ POD modes are defined as
@@
where $i$ and $j$ are paired POD modes, and $N_{\phi}$ is the number of instantaneous velocity fields used in the phase average at corresponding phase angle $\phi$. The mean velocity is combined with the phase-averaged velocity fluctuations to obtain the phase-averaged velocity
@@
from which the phase-averaged out-of-plane vorticity is computed as
@@
The phase-averaged planar Reynolds stress is obtained by averaging the product of the streamwise and transverse components of the instantaneous velocity fluctuations at a given phase angle
$\widetilde{u^{\prime}v^{\prime}}(\boldsymbol{x},\phi)=\frac{1}{N_{\phi}}\sum_{k=1}^{K}\big(a_{k}(t)\psi^{u}_{k}(\boldsymbol{x})\big)\big(a_{k}(t)\psi^{v}_{k}(\boldsymbol{x})\big)|\,\phi_{i,j}=\phi,$ (56)
where $\psi^{u}_{k}(\boldsymbol{x})$ and $\psi^{v}_{k}(\boldsymbol{x})$ are the streamwise and transverse components of the $k^{th}$ POD mode, respectively. The TKE of the phase-averaged velocity is given by
@@
Figure 15: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the first and second POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 9 for an animated version of this figure.
@@
Figure 16: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the first HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 10 for an animated version of this figure.
@@
The first and second POD modes, shown in figure 15, and the first HPOD mode, shown in figure 16, represent a flapping motion in the wake. The phase-averaged velocity obtained using the first and second POD modes and the first HPOD mode, shown in figure 15a and 16a, respectively, are qualitatively indistinguishable, as is the phase-averaged vorticity, shown in figures 15b and 16b, respectively. The structures in the planar Reynolds stresses from the phase-averaged POD modes, shown in figure 15c, are longer and more intense for $x/D<3$ than those in the first HPOD mode, shown in figure 16c, and disappear after $x/D=3$. The structures in the planar Reynolds stress associated with the HPOD mode are shorter and more consistent in size across the measured domain. The intensity of the structures is lower than that of the POD modes, but decays more slowly, with structure remaining visible beyond $x/D=3$. The phase-averaged TKE of the POD modes, shown in figure 15d, appears in periodically shed packets that increase in intensity until $x/D\approx 3$, and then decay after $x/D\approx 3$. The TKE of the HPOD mode, shown in figure 16d, exhibits the same periodic structures as the POD modes but with greater intensity. This is consistent with the TKE of the analytic signal of the turbulent fluctuations being significantly greater than the actual turbulent fluctuations.
@@
Figure 17: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the fourth and fifth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 11 for an animated version of this figure.
@@
Figure 18: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the second HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 12 for an animated version of this figure.
@@
The structure represented by the fourth and fifth POD modes and the second HPOD mode corresponds to a pulsation in the streamwise fluctuations, as evidenced by the deformation of the low-velocity region behind the sphere, shown in figure 17a and 18a, respectively. The streamwise pulsation is also evident in the phase-averaged out-of-plane vorticity, where the high-vorticity regions on either side of the sphere extend further toward the centreline near the sphere, as shown in figures 17b(ii) and 18b(ii). This extension toward the centreline occurs at approximately $x/D=2$, as shown in figures 17b(v) and 18b(v). While the structures in the planar Reynolds stress vary considerably in shape and size in the phase-averaged POD modes, shown in figure 17c, they remain relatively uniform in size and shape in the HPOD mode, as shown in figure 18c. The size and shape of the phase-averaged TKE associated with the POD modes, shown in figure 17d, are more uniform than those of the planar Reynolds stresses. These structures are similar to, but less planar-symmetric than, those in the phase-averaged TKE of the HPOD mode, which also exhibit greater intensity and a more uniform pattern. The intensity of the TKE increases immediately behind the sphere up to $x/D=2$, as seen in figure 18d(ii), before decreasing further downstream, as seen in figure 18d(iv).
@@
Figure 19: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the sixth and eighth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 13 for an animated version of this figure.
@@
Figure 20: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the third HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 14 for an animated version of this figure.
@@
The sixth and eighth POD modes correspond to a flapping motion in the wake, similar to that of the first and second POD modes, but with a shorter wavelength. Consequently, the phase-averaged velocity, shown in figure 19a, and phase-averaged vorticity, shown in figure 19b, closely resemble those of the first and second POD modes. The smaller structures are evident in the phase-averaged planar Reynolds stress, shown in figure 19c, particularly near the sphere. Further downstream of the sphere, the structures increase in streamwise extent, as illustrated in figure 19c(iv). The smaller structures in the phase-averaged TKE, shown in figure 19d, are more distinctly separated across the sphere’s centreline than in the first and second POD modes, where larger structures extend to the centreline. The structures in the phase-averaged planar Reynolds stresses of the third HPOD mode, shown in figure 20c, exhibit more uniformity in size, shape and arrangement compared to those of the POD modes. Additionally, a region with low planar Reynolds stress occurs around $x/D=2$ for all phase angles shown in figure 20b, corresponding to smaller structures in the streamwise turbulent fluctuations in this region, shown in figure 8b. The structures in the phase-averaged TKE are more distinctly defined for the HPOD modes, as shown in figure 20d.
@@
Figure 21: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the ninth and tenth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 15 for an animated version of this figure.
@@
Figure 22: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the fourth HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 16 for an animated version of this figure.
@@
The structure represented by the ninth and tenth POD modes and the fourth HPOD mode resembles that of the fourth and fifth POD modes and the second HPOD mode, namely a periodic streamwise pulsation in the wake. The phase-averaged velocity and vorticity for the POD modes and HPOD mode, shown in figures 21a–b and 22a–b, respectively, closely resemble those of the lower-order modes. The structures in the planar Reynolds stress phase-averaged using the POD modes, shown in figure 21c, are small and rounded in the near wake of the sphere but become larger and more ovular further downstream, with the latter angled away from the centreline in alignment with the V-shaped structures in the POD modes. While the arrangement of these structures varies considerably near the sphere in the POD mode phase average, the planar Reynolds stress of the HPOD mode, shown in figure 22c, clearly exhibits rounded structures for $x/D<2$ and more V-shaped structures for $x/D>2$. Both the POD modes and HPOD modes exhibit a pronounced reduction in the planar Reynolds stress near $x/D=2$, which separates the rounded structures from the V-shaped ones. However, the HPOD modes display this transition more prominently, whereas the POD modes exhibit additional structures in the very near wake which are absent from the HPOD modes. The transition from round structures near the sphere to V-shaped structures further downstream is evident in the phase-averaged TKE, shown in figure 21d and 22d, for the POD modes and the HPOD mode, respectively. This also indicates that the structures near the sphere are more intense and decay towards $x/D=2$, whereas the V-shaped structures increase in intensity around $x/D=3$ before gradually weakening downstream. Similar to the planar Reynolds stress, the structures in the TKE are more distinctly defined in the HPOD mode due to the reduced significance of non-propagating structures in the analytic signal.
@@
The structural details revealed by the POD mode pairs identified from the HPOD modes differ from those of the corresponding HPOD modes. The primary distinction between these methods stems from emphasis on propagating structures in the leading HPOD modes, whereas non-propagating structures are relegated to higher-order modes. Phase-averaging the resulting modes yields more uniformly shaped structures, thereby facilitating the identification and interpretation of propagating structures. However, the reduced emphasis on non-propagating structures may result in a less accurate representation of flow features that do not propagate in the direction of the Hilbert transform, such as those confined to the near wake of the sphere. Although applying the Hilbert transform directly to the POD modes is computationally more efficient than performing HPOD to pair POD modes, it also artificially imposes propagation on structures that may not physically propagate. Thus, both methods of pairing POD modes introduce non-physical assumptions into the flow analysis, which should be carefully considered when interpreting the results.
@@ 5 Concluding Remarks
The velocity in the wake of a sphere at $Re_{D}=7780$ was measured using 2C-2D MCCDPIV, with POD and HPOD applied to the turbulent fluctuations. The resulting POD and HPOD modes were compared to assess whether the HPOD modes, which emphasise propagating structures, could be used to identify POD mode pairs representing the same structure at different phase angles. The analytic signal of the POD modes was then employed as a substitute for the HPOD modes to identify paired POD modes without performing the computationally intensive HPOD. The POD mode pairs identified using these methods were then used to phase-average the turbulent fluctuations to enable the examination of propagating structures in the sphere’s wake.
The leading HPOD modes are shown to represent the same structures as the leading POD modes, with a greater emphasis on the propagating structures in the former. While each HPOD mode consists of a real and imaginary part corresponding to a phase-shift of $\pi/2$ in the mode, this comes at the cost of spectral leakage due to the assumption of periodicity imposed by the Hilbert transform. As a result, it can be difficult to differentiate between genuine flow structures and artefacts introduced by the transform, particularly when taking two-dimensional measurements of a three-dimensional flow. Performing the Hilbert transform on the POD modes to create an analytic signal of the POD modes is shown to identify the same POD mode pairs as the HPOD modes. This method allows the same structures to be identified, without introducing the non-physical elements introduced by the HPOD. The identified POD mode pairs represent flapping and pulsating motions in the wake of the sphere.
As the velocity fields in the present study are not temporally resolved, the Hilbert transform was applied in the streamwise direction, meaning the identified structures are those propagating along this axis. As a result, the temporal information of the structures is not captured by either decomposition. However, the paired POD modes offer insight into the dynamics of turbulent structures, even when flow complexity makes it difficult to identify corresponding POD modes.